# Linguistic Nepotism: Trading-off Quality for Language Preference in Multilingual RAG

### v1

@@ 3.2 Measurement Method
We first construct a citation prediction prompt ending in the form: $x_{i}=s_{i}$[, where the bracket [ signals the start of the citation. To test for language preference for English, we define the set of evaluation languages as $\mathcal{L}_{\mathrm{eval}}=\{\mathrm{en}\}\cup\mathcal{L}_{\mathrm{target}}$, which includes English and all target languages. For each statement, we construct contrastive contexts where only the document to be cited, $d_{c_{i}}$, is presented in a language $\ell\in\mathcal{L}_{\mathrm{eval}}$, while all other evidence documents remain in English. Both $d_{c_{i}}$ and $d_{\neg c_{i}}$ are taken directly from the dataset. The full context is denoted as $\mathrm{Context}(d_{c_{i}}\rightarrow\ell,d_{\neg c_{i}}\rightarrow\mathrm{en})$. Given the prompt prefix $x_{i}$, the model’s next token probability of the correct citation ID token $c_{i}$ corresponding to document $d_{c_{i}}$ conditioned on this context is: $p_{\theta}^{(\ell)}(c_{i})=\mathcal{P}_{\theta}(t=c_{i}|x_{i},q,\mathrm{Context}(d_{c_{i}}\rightarrow\ell,d_{\neg c_{i}}\rightarrow\mathrm{en}))$, where $\mathcal{P}$ is the model’s next token distribution given a prefix, and $\theta$ denotes model parameters. We define the model’s top-predicted citation token as: $\hat{c}^{(\ell)}_{i}=\mathrm{argmax}_{t}(p_{\theta}^{(\ell)})$, and compute citation accuracy in language $\ell$ over $n$ statements as:
@@ 6 Effect of the Query Language
[…]
@@ D.4 Logit Lens Analysis
The model follows a trajectory similar to LLaMA-3.1 8B. Both correct and incorrect citation ID predictions begin to rise around layer 40, peak sharply at layers 52–57, then decline until layer 60 before increasing again and stabilizing toward the final layers. Throughout, correct predictions consistently outnumber incorrect ones. As with LLaMA-3.1 8B, the gap between correct and incorrect predictions narrows for lower-resource languages.
@@ D.5 Query Language Variants
[…]
@@ Appendix E Contributive Attribution Patterns
We conduct the same set of experiments from Section 7 in a setting where the query is in a language other than English. We use the same dataset, MIRACL (231 queries) with machine translated queries, relevant, and irrelevant documents. While fixing the query in target language, we vary the language of one relevant and one irrelevant document under three conditions: (1) tgt-tgt: Both relevant and irrelevant documents are in the target (query) language; (2) tgt-En: Relevant document in the target language, irrelevant document in English; (3) En-tgt: Relevant document in English, irrelevant document in the target language.
@@
As shown in Figure 20, results support the first hypothesis: citation accuracies are generally the lowest when the relevant document is in English and the distractor is in the query language, showing that models preferring language over relevance persists for non-English queries. Interestingly, we show that this trend is most evident for (i) lower-resource languages such as Bengali (bn) and Swahili (sw) and (ii) models that reported to support all tested target languages (Qwen-3 8B and 14B, Gemma-3 27B; Table 4). Conversely, we show mixed results for the second hypothesis. The citation accuracies of tgt-En and tgt-tgt are largely similar, suggesting that English distractors not necessarily easier at misleading models than those in the target language. Overall, our results imply that models show a consistent preference for the query language over relevance, but the distractor’s language matters less when the query is not in English.

### v2

@@ 3.2 Measurement Method
We first construct a citation prediction prompt ending in the form: $x_{i}=s_{i}$[, where the bracket [ signals the start of the citation. To test for language preference for English, we define the set of evaluation languages as $\mathcal{L}_{\mathrm{eval}}=\{\mathrm{en}\}\cup\mathcal{L}_{\mathrm{target}}$, which includes English and all target languages. For each statement, we construct contrastive contexts where only the document to be cited, $d_{c_{i}}$, is presented in a language $\ell\in\mathcal{L}_{\mathrm{eval}}$, while all other evidence documents remain in English. The full context is denoted as $\mathrm{Context}(d_{c_{i}}\rightarrow\ell,d_{\neg c_{i}}\rightarrow\mathrm{en})$. Given the prompt prefix $x_{i}$, the model’s next token probability of the correct citation ID token $c_{i}$ corresponding to document $d_{c_{i}}$ conditioned on this context is: $p_{\theta}^{(\ell)}(c_{i})=\mathcal{P}_{\theta}(t=c_{i}|x_{i},q,\mathrm{Context}(d_{c_{i}}\rightarrow\ell,d_{\neg c_{i}}\rightarrow\mathrm{en}))$, where $\mathcal{P}$ is the model’s next token distribution given a prefix, and $\theta$ denotes model parameters. We define the model’s top-predicted citation token as: $\hat{c}^{(\ell)}_{i}=\mathrm{argmax}_{t}(p_{\theta}^{(\ell)})$, and compute citation accuracy in language $\ell$ over $n$ statements as:
@@ 6 Effect of the Query Language
Setting. We compare the effects of document relevance and language by varying the language of one relevant and one irrelevant document under three conditions: (1) En-En (                                                                                                                                                                                                                                         ): Both relevant and irrelevant documents are in English; (2) tgt-En (): Relevant document in the target language, irrelevant document in English; (3) En-tgt (): Relevant document in English, irrelevant document in the target language. Since ELI5 dataset does not include irrelevant documents, we use MIRACL (Zhang et al., 2023), a multilingual RAG dataset with Wikipedia queries. We use the English subset of the development set, restricting to queries with exactly one relevant document (231 queries). We randomly use one of the irrelevant documents. For each query, we follow the same process described in Section 3. Dataset statistics are in Appendix Table 2.
Figure 5: Accuracy per model with one relevant and one irrelevant evidence document in different languages. : Relevant doc in target language, irrelevant doc in English; : Relevant doc in English, irrelevant doc in target language;  : Baseline, both docs in English. Models trade off document relevance for language preference. Detailed results are in Appendix Figure 5.
@@ 8 Conclusion
Ethics Statement
We recognize that research on language preference in multilingual retrieval-augmented generation can raise issues of fairness, particularly for low-resource and underrepresented languages. Our study aims to highlight such disparities rather than reinforce them. We caution that models may amplify English preference in multilingual contexts, with potential implications for inclusivity and equitable knowledge access.
Reproducibility Statement
All datasets used in our experiments are publicly available (ELI5, MIRACL) and are described in Section 4 and Appendix B, including dataset statistics, splits, and preprocessing steps. Our methodology for constructing multilingual evidence sets, generating English reference reports, and verifying citation correctness is documented in Section 3.1, with prompt templates provided in Appendix A. We report full per-language translation quality scores (Appendix D.1) and include statistical tests to assess the significance of results (Section 3.2, 5.1, and Appendix D). To support reproducibility, we will release our code and processed data upon publication, enabling others to replicate both our measurement pipeline and analyses.
@@ D.5 Query Language Variants
Figure 18: Accuracy per model with one relevant and one irrelevant evidence document in different languages. : Relevant doc in target language, irrelevant doc in English; : Relevant doc in English, irrelevant doc in target language;  : Baseline, both docs in English.
@@ Appendix E Contributive Attribution Patterns
We conduct the same set of experiments from Section 7 in a setting where the query is in a language other than English. We use the same dataset, MIRACL (231 queries) with machine translated queries, relevant, and irrelevant documents. While fixing the query in target language, we vary the language of one relevant and one irrelevant document under three conditions: (1) tgt-tgt (): Both relevant and irrelevant documents are in the target (query) language; (2) tgt-En (): Relevant document in the target language, irrelevant document in English; (3) En-tgt (): Relevant document in English, irrelevant document in the target language.
@@
As shown in Figure 20, results support the first hypothesis: citation accuracies are generally the lowest when the relevant document is in English and the distractor is in the query language, showing that models preferring language over relevance persists for non-English queries. Interestingly, we show that this trend is most evident for (i) lower-resource languages such as Bengali (bn) and Swahili (sw) and (ii) models that reported to support all tested target languages (Qwen-3 8B and 14B, Gemma-3 27B; Table 4). Conversely, we show mixed results for the second hypothesis. The citation accuracies of tgt-En and tgt-tgt are largely similar, suggesting that English distractors are not necessarily easier at misleading models than those in the target language. Overall, our results imply that models show a consistent preference for the query language over relevance, but the distractor’s language matters less when the query is not in English.
@@ Appendix G Alternative Machine Translation System
Appendix H Usage of Large Language Models
We used LLMs to support and refine the writing of our work. Importantly, we did not rely on them to generate content or sentences from scratch. Instead, we employed them primarily to polish the clarity and expression of how we presented our results. In addition, we used them for stylistic adjustments, such as improving readability and removing layout issues (e.g., widows and orphans).